# JEL: C81, O32

### v1

@@ B.2. Benchmark Comparison to Standard Machine Learning Methods
Figure 2 displays receiver operating characteristics for the each class of models estimated with both types of embedding. Error rates are estimated in the testing data. The performance of the receiver operating characteristic for the “ensemble” model whose performance is displayed in Figure 3 is given in purple. The fine-tuned large language model developed in this paper substantially out-paces the performance of standard machine learning methods.
B.3. Prompt Design and Error Analysis
We extract weak labels for 64,000 randomly selected abstracts with two proprietary large language models: OpenAI’s GPT-3.5 and GPT-4 (Nori et al.,, 2023; Bubeck et al.,, 2023). A priori, we expect GPT-4 to output labels of slightly lower quality than hand-labeled output and GPT-3.5 to produce slightly noisier labels than GPT-4. The extraction of GPT-3.5 labels is substantially faster and cheaper.
To produce these labels at scale, each of the two models must be appropriately prompted. That is, each pre-trained model must be provided with a block of text as an input and asked to return the appropriate completion of this input. A priori, it is not clear what prompt structure will work well. The relative infancy of this area of research renders it difficult to identify a set of “best practices.” The proprietary nature of these models makes them difficult to study. See, for details, a guide to prompt engineering from OpenAI here: https://platform.openai.com/docs/guides/prompt-engineering. All of the prompts that we consider are displayed as figures in Appendix D.
We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion. The simplest prompt provides a version of our sample definition, 3.1, and the text of an abstract, and asks the model to return ‘TRUE’ of the abstract satisfies these criteria. Otherwise, it returns false. The text of this prompt is displayed in Figure D.8. We refer to this prompt as Prompt 1.0. A second, more complicated prompt provides the same definition, with a set of examples of publication characteristics that do and do not satisfy the definition. The text of this prompt is displayed in Figure D.14. We refer to this prompt as Prompt 2.0. A third, more complex prompt provides the same definitions and examples, but asks the model to return an explanation of why the record does or does not satisfy these criteria. The text of this prompt is displayed in Figure D.17. We refer to this prompt as Prompt 3.0. We devise initial language for each prompt iteratively, using small numbers of records from our hand-labelled validation data.
@@ B.3. Prompt Design and Error Analysis
We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4. We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis. For each instance in which a model returns a label (TRUE/FALSE) that differs from that in the hand-labelled dataset, we inspect the record. We categorize errors into types and subtypes. To make this concrete, suppose that GPT-3.5, given Prompt 1.0 and an abstract that reports the results of a meta-analysis, erroneously classifies the record as clinical trial, per our definition. In LABEL:tab:Error-Analysis, we flag this error under Prompt 1 for GPT-3.5, as an error of type “meta-analysis.” Within meta-analysis, we assign the error to a sub-type. If the abstract explicitly includes terms such as “meta-analysis,” “literature search,” or “literature review,” we categorize this as an explicit error. If the abstract references that the publication is summarizing existing studies, or searching a database for records and collecting their findings, we categorize this as an implicit error. We devise error types based on our inclusion/exclusion criteria, in 3.1, and select sub-types based on common categories of errors.
This exercise is instructive on three margins. First, it highlights the highest performing prompts. Second, it indicates particular types of errors in categorization—which suggest opportunities for more precise language in a prompt. Third, it draws attention to differences in the performance of the two models. Observe that, in the “Other” error type, we include a subtype called “overly literal interpretation of inclusion criteria.” We primary record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample. Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided. GPT-4 makes two errors. Insight that emerges from this exercise, then, is that different prompt structures may be preferable depending on the model used.
@@
We revise each class of prompt based on these findings. For Prompt 1.0, the simplest true/false prompt, we consider three variants. We refer to these revised prompts as Prompts 1.1, 1.2, and 1.3. The text of these prompts is displayed in Figures D.9, D.10, and D.12. For Prompts 2 and 3, we consider one variant each. We refer to these revised prompts as Prompts 2.1 and 3.1. The text of these prompts is displayed in Figures D.15 and D.18. These changes reflect the differences in performance catalogued in LABEL:tab:Error-Analysis.
Table B.1 reports estimates of the true positive rate and false positive rate for each prompt in both models computed in the validation sample. Observe—for example, with Prompts 1.1, 1.2, and 1.3 queried to GPT-4—that even small changes in the text of a prompt yield substantial differences in performance. We attempt a second round of prompt iteration (Prompt 1.3). We observe that performance deteriorates with even small modification. Thus, we select the two highest performing prompts—one for each proprietary model. We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4. These prompts are used to extract 64,000 labels—TRUE/FALSE flags—using GPT-3.5 and GPT-4. We refer to these as our set of “noisy labels.”
Table B.1. Prompt Performance in Validation Data
Panel A: GPT-3.5
Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
1 0 0.202 0.971
1 0.065 0.949
2 0.049 0.934
3 0.056 0.912
2 0 0.081 0.964
1 0.072 0.964
3 0 0.167 0.971
1 0.068 0.956
Panel B: GPT-4
Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
1 0 0.247 0.788
1 0.172 0.898
2 0.037 0.584
3 0.037 0.489
2 0 0.162 0.876
1 0.176 0.905
3 0 0.248 0.722
1 0.171 0.883
Notes: Table B.1 records the performance of each of eight prompt variants in a sample of 1,000 validation dataset records. Panel A reports performance associated with the proprietary model GPT-3.5. Panel B reports analogous statistics for model GPT-4. Prompt 1 asks the model to return TRUE or FALSE. Prompt 2 asks the model to return TRUE or the name of a specific, excluded category. Prompt 3 asks the model to return TRUE or an explanation of why the record should be excluded. Prompt sub-types correspond to various iterations. Sub-type 0 is the initial version of the prompt. Sub-types 1-3, where applicable, are subsequent iterations. Appendix D records the text of each prompt. We use Prompt 2, Sub-Type 0, to extract weak labels using GPT-3.5, and Prompt 2, Sub-Type 1 to extract weak labels using GPT-4.
Table B.2. Baseline Prompt Error Analysis
@@
Notes: LABEL:tab:Error-Analysis categorizes errors documented in the first round of prompt iteration, using Prompts of sub-type 0. See Table B.1 for details on model performance. For each model error—an instance in which a model returned a classification that deviated from hand-labelled data—we examined the associated record. We categorized errors by Type and Sub-Type. Types are drawn from the exclusion restrictions in 3.1. Sub-types describe consistent characteristics of publications that resulted in such an error. We report counts of errors, by model and prompt, of each Type and Sub-type.
B.4. Fine-tuning
@@
B.5. Performance Assessment
In this appendix, we assess the performance of various fine-tuned language models for classifying whether a publication satisfies the restrictions enumerated in 3.1. Figure B.2 displays estimates of the receiver operating characteristic of several fine-tuned large language models. Here, we consider only the performance of the RoBERTa base model. The top and bottom panels are trained on noisy labels extracted from GPT-3.5 and GPT-4, respectively. We vary the quantity of noisy labels used to train the language model. Additionally, in the bottom panel, we display the performance of a fine-tuned model trained with the 1000 hand-labelled observations in the training dataset. The performance of each language model is measured in the testing dataset.
Figure B.2. Receiver Operating Characteristic by Data Size and Origin
Notes: Figure B.2 displays estimates of the the receiver operating characteristic of several fine-tuned language models. The models differ in terms of the source of their training data. Solid lines display the performance of models trained on noisy labels extracted from either GPT-3.5 or GPT-4. The dotted line displays the performance of a model trained on the hand-labeled data in the training dataset. Estimates of the true positive rate and false positive rate of GPT-3.5 and GPT-4 are indicated with black dots.
The fine-tuned models trained with noisy labels extracted from GPT-4 significantly outperform the models trained with noisy labels extracted from GPT-3.5 or with the hand-labels. There appears to be a threshold, where performance dramatically improves, at around 8,000 training labels. The models displayed in Figure 3 were each obtained by using all 64,000 training labels.
B.6. Final Model: Error Analysis
@@
Appendix C Additional Figures and Further Analyses
@@
This appendix collects additional figures and analyses. Figure C.3 displays series analogous to the series displayed in Figure 4, constructed with the Moderate and Liberal sample restrictions. The results are very similar.
@@
One concern with the comparison made in Figure 4 is that researchers may cite a growing number of publications over time. That is, more papers may cite clinical trials over time because citations are becoming, in some sense, “cheaper.” The relevant clinical literature, then, may not be any larger. To alleviate such concerns, Figure C.4 reports counts, where we scale each citation assigned to a paper by the total number of citations in the originating paper. That is, a paper that receives two citations, each from a paper that cites two papers, will have a weighted citation count of 1.
@@
Figure C.5 displays series analogous to Panel A of Figure 5 with the Moderate and Liberal sample restrictions. Figures C.6 and C.7 display heat maps analogous to Panel B of Figure 5, with the Moderate and Liberal sample restrictions and both 3-Year and 5-Year citation counts. In each case, the results are very similar across specifications.
Figure C.3. Growth in Clinical Research, by Sample Stringency
@@
Notes: Figure C.3 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Each series is reported in terms of the percent change relative to its initial value. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
Figure C.4. Growth in Cost-of-Citation Weighted Clinical Research
@@
Notes: Figure C.4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Here, we report the number of scientific papers using a citation-weighted metric. In particular, we scale each citation assigned to a paper by the total number of citations in the originating paper. Each series is reported in terms of the percent change relative to its initial value. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
Figure C.5. Public Funding and Low-Quality Research, by Sample Stringency
@@
Notes: Figure C.5 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019 in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” See Figure 3 for the true positive and false positive rates of these models. The figure displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal.
@@
Figure C.6. Quality over Time, 3-Year Cites, by Sample Stringency
@@
Notes: Figure C.6 displays heat maps measuring the distribution of three-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
Figure C.7. Quality over Time, 5-Year Cites, by Sample Stringency
@@
Notes: Figure C.7 displays heat maps measuring the distribution of five-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
Appendix D Prompt Repository

### v2

@@ C.2. Benchmark Comparison to Standard Machine Learning Methods
Figure 3.2 displays receiver operating characteristics for the each class of models estimated with both types of embedding. Error rates are estimated in the testing data. The performance of the receiver operating characteristic for the “ensemble” model whose performance is displayed in Figure 3.3 is given in purple. The fine-tuned large language model developed in this paper substantially out-paces the performance of standard machine learning methods.
C.3. Prompt Design and Error Analysis
We extract weak labels for 64,000 randomly selected abstracts with two proprietary large language models: OpenAI’s GPT-3.5 and GPT-4 (Nori et al.,, 2023; Bubeck et al.,, 2023). As discussed in the main text, we identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion. We refer to these prompts as Prompt 1.0, Prompt 2.0, and Prompt 3.0, respectively. The text of these prompts are displayed in Figure E.1, Figure E.7, and Figure E.10, respectively.
@@ C.3. Prompt Design and Error Analysis
We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4. We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis. To make this concrete, suppose that GPT-3.5, given Prompt 1.0 and an abstract that reports the results of a meta-analysis, erroneously classifies the record as clinical trial, per our definition. In LABEL:tab:Error-Analysis, we flag this error under Prompt 1 for GPT-3.5, as an error of type “meta-analysis.” Within meta-analysis, we assign the error to a sub-type. If the abstract explicitly includes terms such as “meta-analysis,” “literature search,” or “literature review,” we categorize this as an explicit error. If the abstract references that the publication is summarizing existing studies, or searching a database for records and collecting their findings, we categorize this as an implicit error. We devise error types based on our inclusion/exclusion criteria, in 3.1, and select sub-types based on common categories of errors.
@@
We revise each class of prompt based on these findings. For Prompt 1.0, the simplest true/false prompt, we consider three variants. We refer to these revised prompts as Prompts 1.1, 1.2, and 1.3. The text of these prompts is displayed in Figures E.2, E.3, and E.5. For Prompts 2 and 3, we consider one variant each. We refer to these revised prompts as Prompts 2.1 and 3.1. The text of these prompts is displayed in Figures E.8 and E.11. These changes reflect the differences in performance catalogued in LABEL:tab:Error-Analysis.
Table C.1. Baseline Prompt Error Analysis
@@
Notes: LABEL:tab:Error-Analysis categorizes errors documented in the first round of prompt iteration, using Prompts of sub-type 0. See Table 3.1 for details on model performance. For each model error—an instance in which a model returned a classification that deviated from hand-labelled data—we examined the associated record. We categorized errors by Type and Sub-Type. Types are drawn from the exclusion restrictions in 3.1. Sub-types describe consistent characteristics of publications that resulted in such an error. We report counts of errors, by model and prompt, of each Type and Sub-type.
C.4. Fine-tuning
@@
C.5. Performance Assessment
In this appendix, we assess the performance of various fine-tuned language models for classifying whether a publication satisfies the restrictions enumerated in 3.1. Figure C.2 displays estimates of the receiver operating characteristic of several fine-tuned large language models. Here, we consider only the performance of the RoBERTa base model. The top and bottom panels are trained on noisy labels extracted from GPT-3.5 and GPT-4, respectively. We vary the quantity of noisy labels used to train the language model. Additionally, in the bottom panel, we display the performance of a fine-tuned model trained with the 1000 hand-labelled observations in the training dataset. The performance of each language model is measured in the testing dataset.
Figure C.2. Receiver Operating Characteristic by Data Size and Origin
Notes: Figure C.2 displays estimates of the the receiver operating characteristic of several fine-tuned language models. The models differ in terms of the source of their training data. Solid lines display the performance of models trained on noisy labels extracted from either GPT-3.5 or GPT-4. The dotted line displays the performance of a model trained on the hand-labeled data in the training dataset. Estimates of the true positive rate and false positive rate of GPT-3.5 and GPT-4 are indicated with black dots.
The fine-tuned models trained with noisy labels extracted from GPT-4 significantly outperform the models trained with noisy labels extracted from GPT-3.5 or with the hand-labels. There appears to be a threshold, where performance dramatically improves, at around 8,000 training labels. The models displayed in Figure 3.3 were each obtained by using all 64,000 training labels.
C.6. Final Model: Error Analysis
@@
Appendix D Additional Figures and Further Analyses
This appendix collects additional figures and analyses.
D.1. Robustness of Trends
@@
Figure D.1 displays series analogous to the series displayed in Figure 4.1, constructed with the Moderate and Liberal sample restrictions. The results are very similar.
@@
One concern with the comparison made in Figure 4.1 is that researchers may cite a growing number of publications over time. That is, more papers may cite clinical trials over time because citations are becoming, in some sense, “cheaper.” The relevant clinical literature, then, may not be any larger. To alleviate such concerns, Figure D.2 reports counts, where we scale each citation assigned to a paper by the total number of citations in the originating paper. That is, a paper that receives two citations, each from a paper that cites two papers, will have a weighted citation count of 1.
Figure D.3 displays series analogous to Figure 4.2 with the Moderate and Liberal sample restrictions.
Figure D.4 gives a displays analogous Figure 4.3 for the distribution of citations received by clinical trials in each calendar year originating from all publications. Figure D.5 gives a displays analogous to Figure 4.5 for the distribution of citations received by publications citing clinical trials (three year sample) originating from all publications.
@@
Figures D.6 and D.7 display heat maps analogous to Figure 4.3 and Figure D.4, with the Moderate and Liberal sample restrictions and both 3-Year and 5-Year citation counts. In each case, the results are very similar across specifications.
Figure D.1. Growth in Medical Research, by Sample Stringency
@@
Notes: Figure D.1 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Each series is reported in terms of the percent change relative to its initial value. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
Figure D.2. Growth in Cost-of-Citation Weighted Medical Research
@@
Notes: Figure D.2 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Here, we report the number of scientific papers using a citation-weighted metric. In particular, we scale each citation assigned to a paper by the total number of citations in the originating paper. Each series is reported in terms of the percent change relative to its initial value. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
Figure D.3. Quality Measures, by Sample Stringency
@@
Notes: Figure D.3 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019 in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” See Figure 3.3 for the true positive and false positive rates of these models. The figure displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal.
Figure D.4. Citation Distribution Across Time, All Publications to Clinical Trials
Notes: Figure D.4 displays a heat map measuring the distribution of citations received by clinical trials in each calendar year from all publications. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. The sample of published clinical trials is constructed with the conservative model.
Figure D.5. Citation Distribution Across Time, All Publications to Citing Publications
Notes: Figure D.4 displays a heat map measuring the distribution of citations received by papers that cite clinical trials, in each calendar year, from all publications. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. The sample of published clinical trials is constructed with the conservative model. The sample of papers that cite clinical trials is constructed using a three-year window.
@@
Figure D.6. Distribution of 3-Year Cites, Clinical Trials, by Sample Stringency
@@
Notes: Figure D.6 displays heat maps measuring the distribution of three-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
Figure D.7. Distribution of 5-Year Cites, Clinical Trials, by Sample Stringency
@@
Notes: Figure D.7 displays heat maps measuring the distribution of five-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
D.2. Growth in Citing Papers by Geography
In Section 4.3.2, we document differential trends in the production of publications by type. Specifically, we find essentially stable geographic patterns in the production of clinical trials, whereas we find that large increases in the quantity of papers in our citing sample are driven by research produced outside of the United States. Here, we provide additional details on this geographic fact. In Section A.4, we provide details on the data that are used to construct these plots.
There are two especially striking aspects of Figure 4.4: the quantity of citing papers published by researchers in China is high and sharply rising, and the quantity of (pooled) citing papers from what we refer to as the “rest of the world” set is also high and rising. Table D.1 breaks down the publication of citing papers by country. We elect to focus on countries that produced at least 200 citing papers in 2019—a cut that captures 93 percent of the citing papers in our data. That is, nearly all of the papers in the citing sample are produced by this relatively small number of countries.
We highlight several patterns in Table D.1. First—consistent with its presentation as a stand-alone series in Figure 4.4—the quantity of citing papers produced by Chinese researchers, the change in the size of this set, and its growth over our sample period are distinctive. While the United States continues the produce the largest number of papers in the citing sample, the growth of this set ranks last among the countries listed in Table D.1. Whereas there is 11 percent change in the quantity of citing papers produced by United States-based researchers, there is more a more than 200 percent change for Chinese papers.
Table D.1. Changes in the Number of Papers Citing Clinical Trials by Country
Country Published, 2013 Published, 2019 Change, 2013 to 2019 % Change, 2013 to 2019
China 1724 5620 3896 226.0
Poland 158 311 153 96.8
Japan 1167 2159 992 85.0
Spain 487 892 405 83.2
Italy 1241 2205 964 77.7
Austria 191 319 128 67.0
Korea 635 1045 410 64.6
Netherlands 617 1011 394 63.9
Belgium 211 335 124 58.8
France 695 1075 380 54.7
Denmark 223 343 120 53.8
Austria 597 907 310 51.9
Czech Republic 320 485 165 51.6
Brazil 286 431 145 50.7
Greece 158 228 70 44.3
Sweden 276 391 115 41.7
Germany 1555 2148 593 38.1
Canada 791 1044 253 32.0
India 302 393 91 30.1
Iran 167 215 48 28.7
Great Britain 1680 2104 424 25.2
Turkey 231 259 28 12.1
U.S.A. 11207 12536 1329 11.9
Notes: Table D.1 records the number of papers that cite clinical trials published in the preceding three years in 2013 and 2019, by country. We include all countries that publish more than 200 papers that cite clinical trials in 2019. The fourth column displays the change in the number of ‘citing’ papers between 2013 and 2019. The fifth column displays the percent change in the number of ‘citing’ papers between 2013 and 2019.
D.3. Quantifying Meta-Analyses
In Section 4.3.1, we present three methods that allow us to identify “review” publications—publications that summarize the findings of existing studies—in our set of citing papers. Each approach is imperfect (for reasons that mirror the motivation for the technical contribution of this paper). Here, we provide additional details on these three approaches.
D.3.1. Hand-Labeling Meta-Analyses
From the set of publications that cite clinical trials (three-year sample), we drew 100 observations randomly from each publication year. We inspected records in the first publication year (2013) and last (2019). Each record was reviewed twice. During the first round of review, we collected notes about the content of the publication. We standardized these notes and observed that these notes could be coded into seven distinct categories. We reviewed the records a second time and assigned these standardized codes. Table D.2 reports the composition of these 100-record samples.
The three largest categories in both samples are pre-clinical studies, observational studies, and meta-analyses. We estimate a decrease in the proportion of pre-clinical studies and an increase in the proportion of meta-analyses. By contrast, the proportion of observation studies appears stable.
Table D.2. Composition of Publications that Cite Clinical Trials
Published in 2013 Published in 2019
Clinical Trial 1 0
Editorial 3 2
Pre-Clinical 26 17
Non-Drug RCT 1 4
Protocol 1 1
Case Study 5 4
Observational 31 30
Meta-Analysis 32 41
Notes:Table D.2 records the results of an exercise aimed at characterizing the composition of the set of publications that cite clinical trials (three-year sample). We randomly draw 100 publications from this sample that were published in 2013 and 2019, respectively. We categorize each publication into one of seven categories. We report the number of publications, in each year, that fall into each category.
D.3.2. Text Searches for Meta-Analyses
Recall that, in the process of constructing the training data that were used to fine-tune our large language models, we identified many records in PubMed that were meta-analyses, literature reviews, etc., based on the contents of their text. We reviewed all hand-labeled records from this training process and identified a set of keywords and phrases that appeared most frequently. We constructed a list of phrases and keywords that consistently indicated, to a human labeler, that a record is a meta-analysis or review article. This list is reproduced in Footnote 51.
Using this list of search terms, we search the abstracts of all records in our citing sample. We flag any record with any of the keywords listed in Footnote 51 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 51 as well.
The objective of this exercise is to yield a conservative estimate of the change in the quantity of meta-analyses over our period of interest. In Figure D.8, we plot this trend. Panel A plots the raw count of meta-analyses flagged by this procedure. Panel B plots the proportion of records in our citing sample that are designated as meta-analyses via this search. As expected, this approach captures a small number of records: in 2013, there are 1,606 records in our citing sample flagged as meta-analyses. This rises to 3,372 records in 2019. Although the baseline level this small, this corresponds to a 109 percent increase over this period.
Figure D.8. Trends in Review Articles and Meta-Analyses
Panel A: Counts Panel B: Proportions
Refer to caption Refer to caption
Notes: Figure D.8 displays the results of two exercises aimed at characterizing the change in the quantity of meta-analyses in the set of publications that cite clinical trials (three-year sample) over time. Estimates associated with the keyword search discussed in Section D.3.2 are displayed in purple. Estimates associated with the ‘NLM’ tag query discussed in Section D.3.3 are displayed in purple. Panel A displays counts of the number of publications that satisfy both sets of criteria. Panel B displays proportions of publications that satisfy both sets of criteria.
D.3.3. NLM Tag Queries for Meta-Analyses
Next, we observe that NLM tags, added when publication records are indexed in PubMed, also record information that may signal whether a record is a meta-analysis. On inspection, three of the 77 NLM tags are relevant: “literature review,” “meta-analysis,” and “systematic review.” Of course, as Section 3.2 makes clear, NLM tags may overcount records. As Section D.3.2 generates what is likely a lower bound, we proceed with this strategy, noting that it is likely to yield an upper bound.
We flag all records in our citing sample indexed with one of these three NLM tags. We plot the resulting trend—as a count in Panel A and as a proportion of records in our citing sample in Panel B—of Figure D.8. In 2013, there are 9,407 records satisfying these criteria. In 2019, there are 14,991 such records. Over our period of interest, there is, thus, a 59 percent increase.
Appendix E Prompt Repository